# Context-aware knowledge graph framework for traffic speed forecasting using graph neural network

### v1

@@ Abstract
Human mobility is intricately influenced by urban contexts spatially and temporally, constituting essential domain knowledge in understanding traffic systems. While existing traffic forecasting models primarily rely on raw traffic data and advanced deep learning techniques, incorporating contextual information remains underexplored due to the lack of effective integration frameworks and the complexity of urban contexts. This study proposes a novel context-aware knowledge graph (CKG) framework to enhance traffic speed forecasting by effectively modeling spatial and temporal contexts. Employing a relation-dependent integration strategy, the framework generates context-aware representations from the spatial and temporal units of CKG to capture spatio-temporal dependencies of urban contexts. A CKG-GNN model, combining the CKG, dual-view multi-head self-attention (MHSA), and graph neural network (GNN), is then designed to predict traffic speed utilizing these context-aware representations. Our experiments demonstrate that CKG’s configuration significantly influences embedding performance, with ComplEx and KG2E emerging as optimal for embedding spatial and temporal units, respectively. The CKG-GNN model surpasses benchmark models, achieving an average MAE of 3.46±0.01 and a MAPE of 14.76±0.09% for traffic speed predictions from 10 to 120 minutes. The dual-view MHSA analysis reveals the crucial role of relation-dependent features from the context-based view and the model’s ability to prioritize recent time slots in prediction from the sequence-based view. The CKG framework’s model-agnostic nature suggests its potential applicability in various applications of intelligent transportation systems. Overall, this study underscores the importance of incorporating domain-specific contexts into traffic forecasting and merging context-aware knowledge graphs with neural networks to enhance accuracy.
@@
Traffic forecasting; Context knowledge graph; Spatial and temporal context; Urban transportation; Graph neural network.
@@ I Introduction
Urban transportation is significantly influenced by surrounding environments due to the supply-demand dynamics of human mobility in the urban space [1, 2]. This environment-related influence manifests across both spatial and temporal contexts, which can benefit the predictive models for traffic speed [3, 4]. Actually, context awareness has emerged as a valuable tool in intelligent transportation systems (ITS) to adapt to various traffic scenarios by fully utilizing multi-source context datasets [5, 6]. Despite its potential on ITS, developing context-aware models to boost traffic forecast and investigate contextual effects remains under exploration.
Short-term traffic state prediction involves forecasting upcoming traffic conditions on road segments for periods within hours [7]. Accurate and efficient predictions are key tasks for successful ITS implementation and support multiple downstream tasks, such as traffic signal optimization. Numerous elegant approaches have been contributed to traffic state prediction research. They can be generally classified into two categories: deductive models based on traffic theory and inductive models based on data analysis [8]. Based on a well-established theoretical background, the models based on traffic theory aim at developing simulation models to mimic traffic behavior. These models are computationally intensive, making it challenging to use them for real-time estimation or short-term forecasting. On the other hand, data-driven models produce predictions by extracting and mapping relationships in traffic data, using techniques ranging from basic time series analysis to deep learning. A thorough examination of these research endeavors can be found in [9, 10]. Nowadays, there is a significant increase in the availability of not only traffic datasets but also context datasets, which provide information about the surrounding environments of traffic systems. This proliferation of context datasets opens new opportunities to further improve traffic forecasting [3].
@@
Despite the potential benefits of using surrounding information in traffic forecasting, the context datasets carrying related knowledge are diverse and complex, deepening the difficulty in integrating them into the prediction task [11]. Generally, urban contexts influencing the traffic system encompass various spatial and temporal factors with a broad scope. From a spatial perspective, the morphological layout of a city impacts how transportation systems are designed and operated [2, 12], which in turn affects human mobility and traffic patterns. Spatial datasets, such as points of interest (POIs), land uses, and transportation networks, are typical examples that describe the spatial dimension of urban contexts [4, 13]. Temporally, the traffic patterns vary significantly throughout time periods, weather conditions, and traffic status [7]. For example, peak hours, typically during morning and evening commutes, require a transportation system capable of handling a high volume of passengers. Conversely, during off-peak hours, the system still needs to operate efficiently but with potentially less demand. Regardless of multiple sources, modalities, and dimensions, these spatio-temporal datasets are vital in shaping the nature of urban transportation, becoming a pivotal role in traffic speed prediction [3]. Therefore, effectively organizing context datasets is the prerequisite to incorporating them into the prediction task.
@@ II-A Knowledge graphs in transportation
Knowledge graphs refer to multi-relational graphs linking various entities through relations, which can structure complex datasets to represent and manage knowledge across domains effectively [18, 19]. These graph-based structures enable the modeling of domain-specific knowledge by organizing datasets into a network of interconnected entities and their relations, thus facilitating comprehensive semantic representations [20]. Hence, knowledge graphs are invaluable in applications ranging from ITS to smart city initiatives, where they contribute significantly to the modeling of urban dynamics and environments [21, 15]. For example, projects like the CARES program have demonstrated the potential of integrating dynamic geospatial knowledge graphs with semantic 3D city databases and agent-based systems for intelligent automation [22].
@@
Essentially, KGE techniques map the entities and relations of a knowledge graph into a multidimensional vector space to obtain their vector representations [17, 26]. This process involves defining a scoring function for each fact to ascertain its plausibility, i.e., a triple of (head entity, relation, tail entity), allowing for the quantitative assessment of the relations and entities within the knowledge graph [19]. Wang et al. [20] categorized KGE approaches into two primary types, i.e., translational distance models (distance-based) and semantic matching models (similarity-based). Translational distance models, including TransE and TransR, evaluate the plausibility of facts by measuring the distance between entities after being translated by the relation vector [18]. An adaptation, KG2E, takes uncertainty into account by modeling entities and relations as vectors from multivariate Gaussian distributions [27]. Semantic matching models, on the other hand, rely on similarity-based scoring functions. They determine the plausibility of facts by matching the latent semantics of entities and relations in their vector space embeddings. Examples of this approach include RESCAL, ComplEx, and NTN, which leverage the interplay of vector space dimensions to encode relational patterns [20]. Additionally, several advanced variants of KGE techniques incorporate extra information into the embeddings, such as entity types, relation paths, textual descriptions, and logical rules [18, 20]. These kinds of variants can enrich the representational capacity of knowledge graphs by considering their spatio-temporal dependencies. In this study, we aim to design a context-aware knowledge graph for traffic speed forecasting considering the spatio-temporal relations of available context datasets. Furthermore, we develop a relation-dependent integration strategy to capture the spatio-temporal path dependencies of urban contexts.
@@ II-B Deep learning in traffic forecasting
So far, different variants of deep learning models have emerged to forecast traffic speed, including convolutional neural networks (CNN), recurrent neural networks (RNN), and graph neural networks (GNN). CNN-based methods are pivotal for grid-based prediction tasks by leveraging their proficiency in processing spatial data and analyzing traffic patterns within grid-like structures [10]. However, applying CNN-based models to traffic forecasting is constrained due to the graph structure of transportation networks. Meanwhile, the inconvenience of modeling temporal dependencies in sequences also severely limits its performance in the prediction task [4]. In contrast, RNN-based models, such as long short-term memory (LSTM) units and gated recurrent units (GRU), have been proposed to capture the temporal dependencies inherent in traffic data, a feat less attainable by conventional machine learning techniques [10]. Despite their potential in modeling time-series data, RNN-based models fail to elucidate the spatial dependencies in transportation networks since traffic status in a given road is significantly affected by its neighboring roads [4]. The development of GNN-based models represents the forefront of efforts to address the need for capturing both spatial and temporal dependencies in traffic forecasting [7, 28]. Variants such as the diffusion convolution recurrent neural network (DCRNN) [29], spatial-temporal graph convolutional network (STGCN) [30], and temporal graph convolutional network (TGCN) [31] are some typical examples of utilizing GNNs effectively for traffic forecasting. These models underpin the current study’s approach to predicting traffic status, showcasing the advanced capabilities of GNNs in this domain.
@@
Furthermore, urban contexts are also important for traffic forecasting by influencing and shaping human mobility patterns [32, 13]. Tedjopurnomo et al. [11] have underscored the critical role of context information in enhancing traffic forecasting, highlighting the benefits of integrating context-aware insights into predictive models. Nowadays, context-aware modeling is increasingly gaining prominence in the field of traffic forecasting. For example, Zhu et al. [33] proposed an attribute-augmented spatio-temporal graph convolutional network (AST-GCN) to model external factors as dynamic and static attributes for traffic prediction. Zhang et al. [3] specified multimodal contexts from spatial and temporal views, and then proposed a multimodal context-based graph convolutional neural network (MCGCN) model to fuse context datasets into traffic prediction. These examples have verified the superiority of using context information in traffic forecasting compared to existing GNN-based models. Generally, the integration of context-aware strategies into deep learning models represents a pivotal development in traffic forecasting, which lays a robust foundation for this study’s focus on speed prediction by leveraging knowledge graphs.
@@ III-A Problem statement
Assuming $V$, $S$, and $T$ represent raw traffic speed at road segments, spatial context datasets, and temporal context datasets, respectively, the traffic forecast problem is defined as predicting traffic speed in the next $b$ time slots utilizing $V$, $S$, and $T$ in the past $a$ time slots. To solve this problem, this study proposes a CKG framework for traffic forecasting using GNN, i.e., a CKG-GNN, to construct context-aware knowledge graphs and then incorporate the context-aware representations into traffic speed prediction. As shown in Figure 1, the proposed CKG-GNN model consists of three parts: (1) spatio-temporal knowledge graph construction, (2) relation-dependent knowledge graph embedding and integration, and (3) context-aware traffic forecast using dual-view MHSA.
Figure 1: An overview of the CKG-GNN model for traffic speed forecasting.
@@ III-B2 Temporal unit construction
Meanwhile, the structure of $\mathbb{K}_{T}$ remains constant over time, while its attributes evolve due to the dynamics of temporal contexts. This setting avoids the need for re-embedding entities and relations within $\mathbb{K}_{T}$ at each time step in the prediction task, thereby improving efficiency while still capturing the essential temporal dynamics.
@@ III-C1 Knowledge graph embedding
KGE techniques are divided into two categories based on their scoring functions: distance-based and similarity-based models [20]. Distance-based models measure the plausibility of facts within $\mathbb{K}$ by calculating the distance between entities following a relation-specific translation. On the other hand, similarity-based models measure the plausibility by matching the latent semantics of entities and relations within their vector spaces. In our study, we delve into both model types, examining their efficacy in CKG embedding and choosing the most suitable methods for embedding $\mathbb{K}_{S}$ and $\mathbb{K}_{T}$.
@@
To compare the difference of embedding $\mathbb{K}$ using different methods, we evaluate their performance on link prediction tasks using the PyKEEN library, where the goal is to predict missing relationships in a fact $(h,r,t)$ [34, 26]. The evaluation is distinguished by the side of the prediction, i.e., (1) right-side: predicting $t$ using $h$ and $r$, (2) left-side: predicting $h$ using $r$ and $t$, and (3) both-side: combining both right-side and left-side evaluation. Here, the mean rank (MR) is used to assess the performance of embedding $\mathbb{K}_{S}$ and $\mathbb{K}_{T}$, which averages the ranks assigned to the correct entities within a prediction list [26]. Its value range is $[1,\infty)$, with lower MR values signifying more accurate embeddings. When ranking the prediction list for MR computation, we employ the realistic rank representing the expected value over all permutations respecting the sort order [34].

### v2

@@ Abstract
Human mobility is intricately influenced by urban contexts spatially and temporally, constituting essential domain knowledge in understanding traffic systems. While existing traffic forecasting models primarily rely on raw traffic data and advanced deep learning techniques, incorporating contextual information remains underexplored due to insufficient integration frameworks and the complexity of urban contexts. This study proposes a novel context-aware knowledge graph (CKG) framework to enhance traffic speed forecasting by effectively modeling spatial and temporal contexts. Employing a relation-dependent integration strategy, the framework generates context-aware representations from the spatial and temporal units of CKG to capture spatio-temporal dependencies of urban contexts. A CKG-GNN model, combining the CKG, dual-view multi-head self-attention (MHSA), and graph neural network (GNN), is then designed to predict traffic speed utilizing these context-aware representations. Our experiments demonstrate that CKG’s configuration significantly influences embedding performance, with ComplEx and KG2E emerging as optimal for embedding spatial and temporal units, respectively. The CKG-GNN model establishes a benchmark for 10-120 min predictions, achieving average MAE, MAPE, and RMSE of 3.46±0.01, 14.76±0.09%, and 5.08±0.01, respectively. Compared to the baseline DCRNN model, integrating the spatial unit improves the MAE by 0.04 and the temporal unit by 0.13, while integrating both units further reduces it by 0.18. The dual-view MHSA analysis reveals the crucial role of relation-dependent features from the context-based view and the model’s ability to prioritize recent time slots in prediction from the sequence-based view. Overall, this study underscores the importance of merging context-aware knowledge graphs with graph neural networks to improve traffic forecasting.
@@
Traffic forecasting; Context-aware knowledge graph; Spatial and temporal context; Urban transportation; Graph neural network.
@@ I Introduction
Urban transportation is significantly influenced by surrounding environments due to the supply-demand dynamics of human mobility in the urban space [1]. This environment-related influence manifests across both spatial and temporal contexts, which can benefit the predictive models for traffic speed [2, 3]. Actually, context awareness has emerged as a valuable tool in intelligent transportation systems (ITS) to adapt to various traffic scenarios by fully utilizing multi-source context datasets [4]. Despite its potential on ITS, developing context-aware models to boost traffic forecast and investigate contextual effects remains under exploration.
Short-term traffic state prediction involves forecasting upcoming traffic conditions on road segments for periods within hours [5]. Recent developments have predominantly focused on advanced deep learning methods, with graph neural networks standing out by leveraging spatio-temporal dependencies of traffic data [6]. However, human mobility occurs in situation-dependent settings, significantly influenced by external factors such as land use and transportation networks [7]. These factors constitute the broader context that permeates urban environments, shaping and constraining individual movements across road segments [8]. For example, points of interest (POIs) and land use data illustrate the spatial arrangement of critical locations and facilities, enhancing our understanding of the environment surrounding transportation networks; traffic jams and weather data provide real-time insights into congestion and meteorological events, respectively, aiding in traffic prediction under various conditions. Despite the critical role of contextual information, existing studies have largely centered on utilizing raw traffic data alongside deep learning techniques for forecasting tasks [3], neglecting the rich contextual information that could enhance predictions in complex urban environments. Therefore, there is significant potential in developing context-based neural-symbolic methods to further improve traffic forecasting by integrating graph neural networks with knowledge representation techniques [2, 9].
@@
Despite the potential benefits of using surrounding information in traffic forecasting, the context datasets carrying related knowledge are diverse and complex, deepening the difficulty in integrating them into the prediction task [6]. Generally, urban contexts influencing the traffic system encompass various spatial and temporal factors with a broad scope. From a spatial perspective, the morphological layout of a city impacts how transportation systems are designed and operated [1, 10], which in turn affects human mobility and traffic patterns. Spatial datasets, such as POIs, land uses, and transportation networks, are typical examples that describe the spatial dimension of urban contexts [3, 8]. Temporally, the traffic patterns vary significantly throughout time periods, weather conditions, and traffic status [5]. For example, peak hours, typically during morning and evening commutes, require a transportation system capable of handling a high volume of passengers. Conversely, during off-peak hours, the system still needs to operate efficiently but with potentially less demand. Regardless of multiple sources, modalities, and dimensions, these spatio-temporal datasets are vital in shaping the nature of urban transportation, becoming a pivotal role in traffic speed prediction [2]. Therefore, effectively organizing context datasets is the prerequisite to incorporating them into the prediction task.
@@ II-A Knowledge graphs in transportation
Knowledge graphs refer to multi-relational graphs linking various entities through relations, which can structure complex datasets to represent and manage knowledge across domains effectively [15, 16]. These graph-based structures enable the modeling of domain-specific knowledge by organizing datasets into a network of interconnected entities and their relations, thus facilitating comprehensive semantic representations [17]. Hence, knowledge graphs are invaluable in applications ranging from ITS to smart city initiatives, where they contribute significantly to the modeling of urban dynamics and environments [18]. For example, projects like the CARES program have demonstrated the potential of integrating dynamic geospatial knowledge graphs with semantic 3D city databases and agent-based systems for intelligent automation [19].
@@
Essentially, KGE techniques transform discrete graph structures into a continuous vector space, where entities and relations within a knowledge graph are encoded as vector representations [14, 22], thereby facilitating their integration into traffic forecasting tasks. This process involves defining a scoring function for each fact to ascertain its plausibility, i.e., a triple of (head entity, relation, tail entity), allowing for the quantitative assessment of the relations and entities within the knowledge graph [16]. Wang et al. [17] categorized KGE approaches into two primary types, i.e., translational distance models (distance-based) and semantic matching models (similarity-based). Translational distance models, including TransE and TransR, evaluate the plausibility of facts by measuring the distance between entities after being translated by the relation vector [15]. An adaptation, KG2E, takes uncertainty into account by modeling entities and relations as vectors from multivariate Gaussian distributions [23]. Semantic matching models, on the other hand, rely on similarity-based scoring functions. They determine the plausibility of facts by matching the latent semantics of entities and relations in their vector space embeddings. Examples of this approach include RESCAL, ComplEx, and NTN, which leverage the interplay of vector space dimensions to encode relational patterns [17]. Additionally, several advanced variants of KGE techniques incorporate extra information into the embeddings, such as entity types, relation paths, textual descriptions, and logical rules [15, 17]. These kinds of variants can enrich the representational capacity of knowledge graphs by considering their spatio-temporal dependencies. In this study, we aim to design a context-aware knowledge graph for traffic speed forecasting considering the spatio-temporal relations of available context datasets. Furthermore, we develop a relation-dependent integration strategy to capture the spatio-temporal path dependencies of urban contexts.
@@ II-B Deep learning in traffic forecasting
So far, different variants of deep learning models have emerged to forecast traffic speed, including convolutional neural networks (CNN), recurrent neural networks (RNN), and graph neural networks (GNN). CNN-based methods are pivotal for grid-based prediction tasks by leveraging their proficiency in processing spatial data and analyzing traffic patterns within grid-like structures [24]. However, applying CNN-based models to traffic forecasting is constrained due to the graph structure of transportation networks. Meanwhile, the inconvenience of modeling temporal dependencies in sequences also severely limits its performance in the prediction task [3]. In contrast, RNN-based models, such as long short-term memory (LSTM) units and gated recurrent units (GRU), have been proposed to capture the temporal dependencies inherent in traffic data, a feat less attainable by conventional machine learning techniques [24]. Despite their potential in modeling time-series data, RNN-based models fail to elucidate the spatial dependencies in transportation networks since traffic status in a given road is significantly affected by its neighboring roads [3]. The development of GNN-based models represents the forefront of efforts to address the need for capturing both spatial and temporal dependencies in traffic forecasting [5, 25]. GNNs are neural architectures that process data represented as graphs by aggregating and updating node features through their connections, making them well-suited for tasks related to transportation networks [6]. Variants such as the diffusion convolution recurrent neural network (DCRNN) [26], spatial-temporal graph convolutional network (STGCN) [27], and temporal graph convolutional network (TGCN) [28] are some typical examples of utilizing GNNs effectively for traffic forecasting. These models underpin the current study’s approach to predicting traffic status, showcasing the advanced capabilities of GNNs in this domain.
@@
Furthermore, urban contexts are also important for traffic forecasting by influencing and shaping human mobility patterns [7, 8]. Tedjopurnomo et al. [6] have underscored the critical role of contextual information in enhancing traffic forecasting, highlighting the benefits of integrating context-aware insights into predictive models. Nowadays, context-aware modeling is increasingly gaining prominence in the field of traffic forecasting. For example, Zhu et al. [29] proposed an attribute-augmented spatio-temporal graph convolutional network (AST-GCN) to model external factors as dynamic and static attributes for traffic prediction. Zhang et al. [2] specified multimodal contexts from spatial and temporal views, and then proposed a multimodal context-based graph convolutional neural network (MCGCN) model to fuse context datasets into traffic prediction. The MCGCN model employs two distinct approaches for managing spatial and temporal contexts: hierarchical spatial embedding, which generates spatial representations by organizing spatial contexts across various dimensions, and multivariate temporal modeling, which learns temporal representations by capturing the dependencies within multivariate temporal contexts. These examples have verified the superiority of using contextual information in traffic forecasting compared to existing GNN-based models. Generally, the integration of context-aware strategies into deep learning models represents a pivotal development in traffic forecasting, which lays a robust foundation for this study’s focus on speed prediction by leveraging knowledge graphs. Building upon existing context-aware strategies, our study aims to propose a CKG framework that utilizes knowledge graphs for modeling domain-specific information in the transportation field, and then integrates this framework into GNNs to improve traffic speed forecasting.
@@ III-A Problem statement
Assuming $V$, $S$, and $T$ represent raw traffic speed at road segments, spatial context datasets, and temporal context datasets, respectively, the traffic forecast problem is defined as predicting traffic speed in the next $b$ time slots utilizing $V$, $S$, and $T$ in the past $a$ time slots. To solve this problem, this study proposes a CKG framework for traffic forecasting using GNN, i.e., a CKG-GNN, to construct context-aware knowledge graphs and then incorporate the context-aware representations into traffic speed prediction. As shown in Figure 1, the proposed CKG-GNN model consists of three parts: (1) spatio-temporal knowledge graph construction, (2) relation-dependent knowledge graph embedding and integration, and (3) context-aware traffic forecast using dual-view MHSA. The code and test datasets of CKG-GNN are available at https://github.com/mie-lab/CKG-traffic-forecasting.
Figure 1: An overview of the CKG-GNN model for traffic speed forecasting. It comprises three parts: spatio-temporal KG construction, relation-dependent KG embedding and integration, and context-aware traffic forecast. The model processes inputs of traffic speed, spatial context, and temporal context data from the past $a$ time steps, and outputs traffic speeds for the forthcoming $b$ time steps.
@@ III-B2 Temporal unit construction
Meanwhile, the structure of $\mathbb{K}_{T}$ remains constant over time, while its attributes evolve due to the dynamics of temporal contexts. Moreover, real-time updates, such as sudden traffic jams or extreme weather events, solely influence the attributes without altering the structure of $\mathbb{K}_{T}$. This setting avoids the need for re-embedding entities and relations within $\mathbb{K}_{T}$ at each time step in the prediction task or when including longer time periods, thereby improving efficiency while still capturing the essential temporal dynamics.
@@ III-C1 Knowledge graph embedding
KGE refers to embedding entities and relations of a knowledge graph into vector representations while preserving the inherent structure of the graph, which can be divided into two categories based on their scoring functions: distance-based and similarity-based models [17]. Distance-based models measure the plausibility of facts within $\mathbb{K}$ by calculating the distance between entities following a relation-specific translation. On the other hand, similarity-based models measure the plausibility by matching the latent semantics of entities and relations within their vector spaces. In our study, we delve into both model types, examining their efficacy in CKG embedding and choosing the most suitable methods for embedding $\mathbb{K}_{S}$ and $\mathbb{K}_{T}$. To facilitate the understanding and comparison of distance-based and similarity-based models, Table A1 in the Appendix summarizes the scoring functions and key features of each knowledge graph embedding technique.
@@
To compare the difference of embedding $\mathbb{K}$ using different methods, we evaluate their performance on link prediction tasks using the PyKEEN library, where the goal is to predict missing relationships in a fact $(h,r,t)$ [30, 22]. The evaluation is distinguished by the side of the prediction, i.e., (1) right-side: predicting $t$ using $h$ and $r$, (2) left-side: predicting $h$ using $r$ and $t$, and (3) both-side: combining both right-side and left-side evaluation. In this task, the commonly used evaluation metrics are based on the ranks of correct predictions, including mean rank (MR), mean reciprocal rank (MRR), and Hits@K (H@K) [17, 22, 16]. MR averages the ranks of correct entities within the prediction list; MRR calculates the average reciprocal rank of the first correct prediction; and H@K assesses the frequency of correct predictions within the top K results, such as H@5. Moreover, to enhance evaluation accuracy by compensating for dataset-specific biases and imbalances, we have implemented the adjusted Hits@K (AH@K) metric, which considers the expectation of correct predictions appearing within the top K results [31]. A lower MR indicates better performance, whereas higher values of MRR, H@K, and AH@K signify improved model performance. When ranking the prediction list for computing evaluation metrics, we employ the realistic rank representing the expected value over all permutations respecting the sort order [30].
To elucidate the process of CKG construction and embedding evaluation, we detail the process in Algorithm 1 using inputs of spatial context $S$ and temporal context $T$. The initial steps involve establishing the spatial unit $\mathbb{K}_{S}$ based on spatial facts in Section III-B1, and the temporal unit $\mathbb{K}_{T}$ based on temporal facts in Section III-B2. The framework then progresses to evaluate various KGE techniques to determine the most suitable for both $\mathbb{K}_{S}$ and $\mathbb{K}_{T}$ using evaluation metrics. The process concludes with the embedding results of $S$ and $T$.
Input: spatialContext $S$, temporalContext $T$
Output: embedResults of $S$ and $T$
INIT spatialUnit $\mathbb{K}_{S}$, temporalUnit $\mathbb{K}_{T}$
for spatialEntities in $S$ do
       ADD spatialFacts $\mathbb{K}_{S}(road)$, $\mathbb{K}_{S}(poi)$, $\mathbb{K}_{S}(land)$ to $\mathbb{K}_{S}$
       ADD spatialLinkFact $\mathbb{K}_{S}(link)$ to $\mathbb{K}_{S}$
for temporalEntities in $T$ do
       ADD temporalFacts $\mathbb{K}_{T}(time)$, $\mathbb{K}_{T}(jam)$, $\mathbb{K}_{T}(weather)$ to $\mathbb{K}_{T}$
       ADD temporalLinkFact $\mathbb{K}_{T}(link)$ to $\mathbb{K}_{T}$
for $\mathbb{K}$ in $\mathbb{K}_{S}$, $\mathbb{K}_{T}$} do
       INIT bestMetrics, bestKGEmbed
       for KGEmbed in {TransE, TransR, KG2E, RESCAL, ComplEx, NTN} do
             EMBED $\mathbb{K}$ using KGEmbed
             COMPUTE {MR, MRR, H@K, AH@K}
             if {MR, MRR, H@K, AH@K} OUTPERFORM bestMetrics then
                   bestMetrics = {MR, MRR, H@K, AH@K}
                   bestKGEmbed = KGEmbed
      embedResults = EMBED $\mathbb{K}$ using bestKGEmbed
       return embedResults
Algorithm 1 CKG construction